# Generative AI models enable efficient and physically consistent sea-ice simulations

### v1

@@ 4.3 Surrogate modelling with censored flow matching
$-\mathrm{I}(x_{t+12\,\text{h},k,l}=x^{U}_{k})\left[\log\Phi\Big{(}\frac{\widehat{v}_{\boldsymbol{\theta},k,l}-u_{k,l}}{s_{k}}\Big{)}\right]$
@@
with $\Phi$ as the normal Gaussian cumulative distribution function (CDF), $\mathrm{I}(\cdot)$ as indicator function, and $(\cdot)_{k,l}$ as scalar that is indexed for the $k$-th variable and the $l$-th grid point. Note that the case for the cost function is changed depending on the true state at time $t+12\,\text{h}$, which differs from the common practise of using activation functions for the bounds during training. For variables that are unbounded like the velocities, the lower and upper bound vanish and only the weighted mean-squared error is left as cost function. When the true state is between the physical bounds, the cost function is the commonly used one for flow matching [24, 23], where the prediction target is the true velocity. When the true state is exactly on the lower or upper bound, the prediction is pushed towards exceeding these bounds in negative or positive direction, respectively. Hence, this cost function effectively modifies the role of the predicted velocity: it regresses the true velocity and classifies if the lower or upper bound is reached.
@@
The censored Gaussian distribution can be derived from applying a static thresholding function [43] to the prediction that follows a Gaussian distribution and, hence, properly accounts for physical bounding in that predictions. To apply the neural network’s output as velocity approximation, we have to project the states to $\tau=1$ and apply the thresholding,
@@
$\widehat{\mathbf{z}}_{1\mid\tau}=\mathrm{minimum}\Big{(}\mathrm{maximum}\big{(}\mathbf{z}_{\tau}+\tau\widehat{\mathbf{v}}_{\boldsymbol{\theta}},(\mathbf{\mathbf{x}^{L}}-\mathbf{x}_{t})/\boldsymbol{\sigma}_{\mathrm{clim}}\big{)},(\mathbf{\mathbf{x}^{U}}-\mathbf{x}_{t})/\boldsymbol{\sigma}_{\mathrm{clim}}\Big{)}$ (9)
with $\mathrm{minimum}(\cdot,\cdot)$ and $\mathrm{maximum}(\cdot,\cdot)$ as minimum and maximum function, respectively. Based on this thresholded state, we recover the thresholded velocity by the relation
$\widehat{\mathbf{v}}_{\boldsymbol{\theta}}=\frac{\widehat{\mathbf{z}}_{1\mid\tau}-\mathbf{z}_{\tau}}{1-\tau},$ (10)
@@
which can be used in any integration scheme to solve the ODE that defines the generative flow, Eq. (3). Effectively, we treat the neural network’s output as median approximation for the velocity, which is thresholded if the bounds are exceeded. This way, we can generate data that lies exactly on possible bounds, while being produced by a deterministic velocity approximation.
4.4 Experimental details
In our experiments, we train two different surrogate models: GenSIM and a deterministic baseline model . To make both models comparable, we use similar experimental settings, which we explain in the following.
GenSIM is based on a transformer architecture, as concisely described in Section 1 and further elucidated in Appendix C.1. Since the transformer scales quadratically with the number of grid points/tokens, we use a U-ViT architecture [62, 35, 34] for the deterministic model: a convolutional U-Net coarse-grains the pan-Arctic data, while we use a transformer in the bottleneck with almost the same configuration as for GenSIM. Due to the additional convolutional blocks, the deterministic model is with $32.2\times 10^{6}$ parameters slightly bigger than GenSIM with $28.1\times 10^{6}$ parameters. To counteract the increased tendency to overfit by the deterministic model, we use the model that performs the best over the validation dataset.
The neural networks are trained with Eq. (7) and Eq. (2) for GenSIM and the deterministic model, respectively. The dataset for both models is augmented by distribution augmentation [47, 35], where the data is randomly flipped in horizontal and vertical direction, and randomly rotated by a $90^{\circ}$ angle. The labels, describing which augmentation is activated, are added to the input of the neural networks.
@@ 4.4 Experimental details
GenSIM is trained with pseudo time steps drawn from a uniform distribution, $\tau\in U(0,1)$, while the weighting of the loss function is learned alongside the model [63], effectively reducing the need to tune the training of the generative flow. The samples are randomly averaged with window sizes drawn from $(1,2,4,8)$ with probabilities $(0.5\overline{33},0.2\overline{66},0.1\overline{33},0.0\overline{66})$ and randomly sliced into patches with $80\times 80$ grid points such that each patch contains at least $10\,\%$ unmasked grid points. For the estimation of the loss function, we remove the $8$-pixel border around each patch to obtain a prediction size of $64\times 64$. In total, we randomly extract eight patches per sample for GenSIM’s training.
@@
Both models are trained with mini-batches of $32$ samples and the AdamW optimiser [64, 65] with $\beta_{1}=0.9$ and $\beta_{2}=0.99$ and a weight decay with $\lambda=10^{-3}$. The learning rate is linearly increased from $\gamma_{\mathrm{min}}=10^{-6}$ to $\gamma_{\mathrm{max}}=5\times 10^{-4}$ within the first $5000$ iterations. Afterwards, the learning rate is decayed by a cosine until the end of the training: $1\times 10^{6}$ iterations for GenSIM and $5\times 10^{5}$ iterations for the deterministic model. After training, we retain the model that has the lowest validation loss: GenSIM reaches its optimum at the end, while the deterministic model starts to overfit after around $1.5\times 10^{5}$ iterations. We track the exponential moving average (EMA) of the model’s weights with a decay rate of $\alpha=0.9999$. We train the models on data for 1995199520142014, giving $913$ iterations per training epoch, while we validate on $2015$. The testing period 2016201620182018 is only revealed after the training has been finished and used to produce the results in this study.
@@
We train on two Nvidia RTX A6000 with $48\,\text{GB}$ memory (per-device batch size $16$) and eight Nvidia A100 with $80\,\text{GB}$ memory (per-device batch size $4$). The training of GenSIM needs on the eight Nvidia A100 around $2$ days. The models are implemented with bfloat16 mixed-precision in Python [66] based on PyTorch [67], PyTorch lightning [68], and Hydra [69]. The data and code for the models will be made publicly available upon acceptance.
@@
As sampler for GenSIM, we use a second-order Heun scheme [61] with 2020 steps and a sigmoid-like pseudo-time scheduler as visualised in Fig. 9b. Each forecast with GenSIM hence needs $39$ neural network evaluations, taking around $4.5\,\text{s}$ on a Nvidia RTX A6000.
4.5 Comparison between GenSIM and deterministic model
For the testing period, 2016–2018, we produce weekly forecasts with GenSIM and the deterministic model. Starting from 1st January of each year, we initialise 10-day forecasts on a weekly basis, performing 153 initialisations in total. The forecasts are initialised with the state from neXtSIM-OPA and use forcings from the ERA5 reanalysis. We run an ensemble with 16 members for GenSIM, while we run a single forecast per initialisation for the deterministic model. These forecasts are then compared to the neXtSIM-OPA simulation as truth.
We evaluate the root-mean-squared error (RMSE) between the forecasts and the truth over all ocean grid points that are covered by sea ice in the neXtSIM-OPA simulation (concentration >0.15>0.15). To estimate a normalised RMSE, we normalise the forecast errors by the standard deviation of the climatology. The spread-skill ratio is estimated as ratio between the root-mean-squared standard deviation from GenSIM’s ensemble members and the RMSE of the ensemble mean with a correction factor of $\sqrt{17/16}$ [70].
The energy spectra are estimated over the central Arctic region with $128\times 128$ grid points and averaged in logarithmic space over all initialisation times. To estimate the extent accuracy, the predicted sea-ice concentration is converted into an extent mask (concentration >0.15>0.15). The accuracy is then the agreement between the prediction and the truth over all ocean grid points.
4.6 Sea-ice volume estimation
@@
We normalise all sea-ice volume datasets into a common format: we use the Cartesian EASE-$25$-km mesh from CS2/SMOS as common mesh on which the volumes are estimated. We interpolate the sea-ice thickness and concentration of each dataset with bilinear interpolation to that mesh.
@@
4.7 Scaling analysis of sea-ice deformation rates
Since GenSIM has been trained on data up to 2014, we split the analysis of the sea-ice deformation rates into two parts: in the first part, we track individual features found by RGPS in the GenSIM and neXtSIM simulations for a qualitative analysis. In the second part, we work on the common curvilinear mesh of GenSIM and neXtSIM, using spatial averaging for a quantitative scaling analysis.
The individual features found by RGPS are grouped with a 12-hourly frequency from 2007-02-01 03:00 to 2007-02-10 15:00 with a ± $\pm\,6\,\text{h}$ window. From this grouped dataset, we retain only features that are far from the coast ($\geq 150\,\text{km}$) and are refound three days later. Based on the temporal and spatial difference, we estimate the velocity of the features by the relation $\mathbf{u}_{t}=(\mathbf{x}_{t+3\,\text{d}}-\mathbf{x}_{t})/\Delta t$, where $\Delta t\in(2\,\text{d}+12\,\text{h},3\,\text{d}+12\,\text{h})$ is the true time difference.
Afterwards, we triangulate each group of features, retaining triangles with all angles larger than $20^{\circ}$ and an area between $25\,\text{km}^{2}$ and $400\,\text{km}^{2}$. The velocity gradients, needed to estimate the deformation rates, are estimated based on the found triangles with the contour integral [9], e.g., $u_{x}=\frac{1}{A}\oint udy$ for the $u$-velocity gradient in $x$-direction with $A$ as triangle area. The shear deformation, $\dot{\epsilon}_{\mathrm{shear}}$, divergent deformation, $\dot{\epsilon}_{\mathrm{div}}$, and total deformation rates $\dot{\epsilon}_{\mathrm{tot}}$ are estimated by
$\dot{\epsilon}_{\mathrm{shear}}$ $=\sqrt{(u_{x}-v_{y})^{2}+(u_{y}+v_{x})^{2}},$ (11)
$\dot{\epsilon}_{\mathrm{div}}$ $=u_{x}+v_{y},$ (12)
$\text{and }\dot{\epsilon}_{\mathrm{tot}}$ $=\sqrt{\dot{\epsilon}^{2}_{\mathrm{shear}}+\dot{\epsilon}^{2}_{\mathrm{div}}}.$ (13)
Every 12-hours, we initialise particles at positions of the found features. Based on the simulated velocity, we advect all particles three-days forward in time for the GenSIM and neXtSIM simulation. As advection scheme, we use a simple forward Euler integration with an hourly time step, a bilinear velocity interpolation in space, and linear interpolation in time. The displacement of the particle within the three days determines then the velocity.
To make a quantitative analysis from 2015-01-01 03:00 to 2018-12-31 15:00, we use the central Arctic slice (see also Fig. 6) of GenSIM’s $1/4^{\circ}$ mesh, circumventing issues with land masses. Instead of advecting particles, we calculate the velocity gradients with the predicted velocities, which correspond to six-hourly averages. The velocity gradients are given by the contour integral over the rectangles that are defined by four neighbouring cells. We estimate the deformation rates based on the velocity gradients as before with Eq. (11)–(13).
As similarly used in other studies [6], we coarse-grain the velocity gradients by area-averaging over increasing window size. The coarse-grained velocity then provide us with deformation rates at that coarse-graining scale. At each scale, we track the first three distributional moments of the deformation rates. The scaling analysis in space assumes that the moments $q$ follow a power law, $\big{\langle}\dot{\epsilon}^{q}(L)\big{\rangle}\sim L^{-\beta(q)}$, which depends on the spatial resolution $L$. The scaling coefficient $\beta(q)$ describes how the moments change by coarse-graining. The dependence of the scaling coefficients on the moment gives an estimate how much the localisation of fracturing events are dependent on the magnitude of these events, a positive curvature indicates thereby a multifractal behaviour [9].
4.8 Estimation of sea-ice volume change
@@
$\frac{\partial\mathrm{vol}_{t}}{\partial t}+\nabla\cdot(\mathbf{u}_{t}\,\mathrm{vol}_{t})=S_{t}.$ (14)
@@
To estimate the sea-ice volume change, we use finite differences, $\frac{\partial\mathrm{vol}_{t}}{\partial t}\approx(\mathrm{vol}_{t+12\,\mathrm{h}}-\mathrm{vol}_{t})/12\,\mathrm{h}$. Working on the Arakawa-C mesh, we further approximate the dynamical part with a second‑order centred finite‑volume advection scheme based on the volume and velocities for time $t$ and assume constant dynamics until time $t+12\mathrm{h}$. The source term $S_{t}$ is then given as difference between the true volume change and the change explained by the approximated dynamics.
[…]
Acknowledgements
This study is a contribution to the SASIP project funded under Grant no. G-24-66154 by Schmidt Science–a philanthropic initiative that seeks to improve societal outcomes through the development of emerging science and technologies. TSF, MB, CD, and AF additionally received financial support from INSU/CNRS by the project DeepGeneSIS (PNTS). This work was granted access to the HPC resources of IDRIS under the allocations 2021-AD011013069, 2022-AD011013069R1, 2023-AD011013069R2, and 2024-AD011013069R3 made by GENCI. We would like to thank Lorenzo Zampieri and Joffrey Dumont Le Brazidec (ECMWF), through discussing they helped to focus this study. Additionally, we acknowledge Guillaume Boutin (NERSC) for providing the training dataset and other members from the SASIP project which gave helpful comments along the way. This work contains data modified from Copernicus Climate Change Service information 2024. Neither the European Commission nor ECMWF is responsible for any use that may be made of the Copernicus information or data it contains. CEREA is a member of the Institut Pierre-Simon Laplace (IPSL).
Appendix A RMSE Score cards
Figure 6: The different regions used throughout the manuscript with the following names: 1=central Arctic; 2=Beaufort sea; 3=Chukchi sea; 4=east Siberian sea; 5=Laptev sea; 6=Kara sea; 7=Barents sea; 8=east Greenland sea; 9=Baffin bay, Labrador seas, and gulf of St. Lawrence; 10=Hudson bay; 11=Canadian archipelago. The black hatched rectangle is the region in the central Arctic used to estimate the spectra and fractal scaling. The regions are extracted from NSIDC’s regions on the EASE-Grid 2.0 with a 3.125 km resolution [71].
@@ Appendix A RMSE Score cards
Inspired by ECMWF’s score cards (https://sites.ecmwf.int/ifs/scorecards/, last access: 19th August 2025), the root-mean-squared error is decomposed into averages across seasons and regions. For that decomposition, we use the indexed regions as indicated in Fig. 6. Mesh cells over the Atlantic ocean that mostly cover open water are omitted from the estimation of the root-mean-squared error (RMSE). As previously, the RMSE is estimated only over ice-covered mesh cells (concentration >0.15>0.15). The score cards (Fig. 7 and Fig. 8) show how GenSIM’s ensemble mean improves the predictions compared to the deterministic surrogate model. Note that there is almost no snow during summer (JAS), making its error estimate unreliable in that season.
Figure 7: The score card, depicting the relative root-mean-squared error (RMSE) of GenSIM’s ensemble mean to the RMSE of the deterministic baseline for different regions, variables, and seasons, as estimated over all ice-covered grid points (sit=thickness, sic=concentration, sid=damage, siu=$u$-velocity, siv=$v$-velocity, snt=snow-on-ice thickness). A blue RMSE corresponds to an improvement of GenSIM, while a red one shows a degradation of the forecast. The regions (1–5) are depicted and explained in Fig. 6 and remaining regions are shown in Fig. 8.
Figure 8: Continuation of Fig. 7 for regions (6–11).

### v2

@@ Acknowledgments
$-\mathrm{I}(x_{t+12\,\text{h},k,l}=x^{U}_{k})\left[\log\Phi\Big(\frac{\widehat{v}_{\boldsymbol{\theta},k,l}-u_{k,l}}{\sigma_{k}}\Big)\right],$
@@
with $\Phi$ as the normal Gaussian cumulative distribution function (CDF), $\mathrm{I}(\cdot)$ as indicator function, and $(\cdot)_{k,l}$ as scalar that is indexed for the $k$-th variable and the $l$-th grid point. Note that the case for the cost function is changed depending on the true state at time $t+12\,\text{h}$, which differs from the common practise of using activation functions for the bounds during training (?, ?). For variables that are unbounded like the drift, the lower and upper bound vanish and only the weighted mean-squared error is left as cost function. When the true state is between the physical bounds, the cost function is the commonly used one for flow matching (?, ?), where the prediction target is the true velocity. When the true state is exactly on the lower or upper bound, the prediction is pushed towards an increase in the probability that the bounds are exceeded in negative or positive direction, respectively. Hence, this cost function effectively modifies the role of the predicted velocity: it regresses the true velocity and classifies if the lower or upper bound is reached.
@@
The censored Gaussian distribution can be derived from applying a static thresholding function (?) to the prediction that follows a Gaussian distribution and, hence, properly accounts for physical bounding in that predictions. To apply the neural network’s output as velocity approximation, we have to project the states to $\tau=1$ and into physical space and apply the thresholding,
@@
$\widehat{\mathbf{x}}_{t+12\,\text{h}\mid s}$ $=\mathbf{x}_{t}+\boldsymbol{\sigma}_{\mathrm{clim}}\odot(\mathbf{z}_{s}+(1-s)\widehat{\mathbf{v}}_{\boldsymbol{\theta}})$ (S8)
$\widehat{\mathbf{x}}^{\textrm{thres}}_{t+12\,\text{h}\mid s}$ $=\begin{cases}\mathbf{x}^{L}&\text{if}~~~\widehat{\mathbf{x}}_{t+12\,\text{h}\mid s}<\mathbf{x}^{L}\\ \mathbf{x}^{U}&\text{if}~~~\widehat{\mathbf{x}}_{t+12\,\text{h}\mid s}>\mathbf{x}^{U}\\ \widehat{\mathbf{x}}_{t+12\,\text{h}\mid s}&\text{otherwise}\end{cases}$ (S9)
@@
with $\mathbf{x}^{L}$ and $\mathbf{x}^{U}$ being the lower and upper bound, respectively. Based on this thresholded state, we recover the thresholded velocity by reversing Equation S8. This thresholded velocity can be used in any integration scheme to solve the ODE that defines the generative flow, Equation S2. Effectively, we treat the neural network’s output as median approximation for the velocity, which is thresholded if the bounds are exceeded. This way, we can generate data that lies exactly on possible bounds, while being produced by a deterministic velocity approximation.
Experimental details
GenSIM is based on a transformer architecture with $28.1\times 10^{6}$ parameters, as concisely described in Section Sea-ice modelling with GenSIM and further elucidated in Appendix Transformer architecture. The training scheme is derived in Surrogate modelling with censored flow matching and the loss function for training is Equation S6. The training dataset is augmented by distribution augmentation (?, ?): the data is randomly flipped in horizontal and vertical direction, and a -$90^{\circ}$ rotation is randomly activated. The labels, describing which augmentation is activated, are added to the input of the neural networks.
@@
GenSIM is trained with pseudo time steps drawn from a uniform distribution, $s\sim U(0,1)$, while the weighting of the loss function is learned alongside the model (?), effectively reducing the need to tune the training of the generative flow. The samples are randomly averaged using window sizes of $(1,2,4,8)$ drawn with probabilities $(0.5\overline{33},0.2\overline{66},0.1\overline{33},0.0\overline{66})$, and subsequently cropped into patches with $80\times 80$ grid points, while ensuring that each patch contains at least $10\,\%$ unmasked grid points. For the estimation of the loss function, we remove the $8$-pixel border around each patch to obtain a prediction size of $64\times 64$. In total, we randomly extract eight patches per sample at various averaging scales for GenSIM’s training.
@@
GenSIM is trained with mini-batches of $32$ samples and the AdamW optimiser (?, ?) with $\beta_{1}=0.9$ and $\beta_{2}=0.99$ and a weight decay with $\lambda=10^{-3}$. The learning rate is linearly increased from $\gamma_{\mathrm{min}}=10^{-6}$ to $\gamma_{\mathrm{max}}=5\times 10^{-4}$ within the first $5000$ iterations. Afterwards, the learning rate is decayed by a cosine until the end of the training: $10^{6}$ iterations for GenSIM. As GenSIM is robust against overfitting, we retain the model at the end of the training. We track the exponential moving average (EMA) of the model’s weights with a decay rate of $\alpha=0.9999$. We train the models on data for 1995199520142014, giving $913$ iterations per training epoch, while we validate on $2015$. The testing period 2016201620182018 is only revealed after the training has been finished and used to produce the results in this study.
@@
We train on eight Nvidia A100 with $80\,\text{GB}$ memory (per-device batch size $4$), on which the training needs around $2$ days. The models are implemented with bfloat16 mixed-precision in Python (?) based on PyTorch (?), PyTorch lightning (?), and Hydra (?). The code for the model, the training, and to produce the training data is available at https://github.com/cerea-daml/gensim, while the trained model checkpoints are available at https://huggingface.co/tobifinn/gensim.
@@
For inference, we use two Nvidia RTX A6000 with $48\,\text{GB}$ memory. As sampler for GenSIM, we use a second-order Heun scheme (?) with $20$ steps and a sigmoid-like pseudo-time scheduler as visualised in Fig. S10b. Each forecast with GenSIM hence needs $39$ neural network evaluations, taking around $4.5\,\text{s}$ on an Nvidia RTX A6000.
Sea-ice volume estimation
@@
We normalise all sea-ice volume datasets into a common format: we use the Cartesian EASE-$25$-km mesh from OSISAF as common mesh on which the volumes are estimated. We interpolate the sea-ice thickness and concentration of each dataset with bilinear interpolation to that mesh.
@@
$\frac{\partial\mathrm{vol}_{t}}{\partial t}+\nabla\cdot(\mathbf{u}_{t}\,\mathrm{vol}_{t})=S_{t}.$ (S10)
@@
To estimate the sea-ice volume change, we use finite differences, $\frac{\partial\mathrm{vol}_{t}}{\partial t}\approx(\mathrm{vol}_{t+12\,\mathrm{h}}-\mathrm{vol}_{t})/12\,\mathrm{h}$. Working on the Arakawa-C mesh, we further approximate the dynamical part with a second‑order centred finite‑volume advection scheme based on the volume and drift for time $t$ and assume constant dynamics until time $t+12\mathrm{h}$. The source term $S_{t}$ is then given as difference between the true volume change and the change explained by the approximated dynamics.
Supplementary Text
Short-term prediction quality
We test GenSIM’s short-term forecasting qualities in a perfect model experiment for 2016–2018, where the simulation of neXtSIM-OPA is our truth, against which we evaluate. Initialised from the simulation and driven by ERA5 forcings, we start a 10-day forecast once a week, effectively performing $153$ initialisations. Running an ensemble of $16$ members from the same initial conditions and forcings, we compare the performance against persistence forecasts and against a deterministic deep learning model.
We use a U-ViT architecture (?, ?, ?) for the deterministic model, as the transformer scales quadratically with the number of grid points/tokens: a convolutional U-Net coarse-grains the pan-Arctic data, while we use a transformer in the bottleneck with almost the same configuration as for GenSIM. Due to the additional convolutional blocks, the deterministic model has $32.2\times 10^{6}$ parameters, slightly more than GenSIM that has $28.1\times 10^{6}$ parameters. To counteract the increased tendency to overfit by the deterministic model, we use the model that performs the best over the validation dataset, which is reached after around $1.5\times 10^{5}$ iterations. All other experimental details but the maximum number of iterations ($5\times 10^{5}$ iterations, important for the learning rate decay) are the same as for GenSIM.
For a 12-hour lead time, the deterministic model predicts normalised tendencies, which are scaled by the per-variable climatological standard deviation. The loss function is then given as the mean-squared error between the predicted scaled tendencies and the true scaled tendencies. Each variable is effectively weighted by the inverse of its climatological variations as similarly used for other multivariate deterministic surrogate models (?, ?). During the autoregressive prediction, we enforce the physical bounds such as non-negativity of sea-ice thickness by clipping (?). Hence, this architecture can be seen as a multivariate and larger version of previously used deterministic surrogate models for sea-ice predictions (?, ?).
We evaluate the root-mean-squared error (RMSE) between the forecasts and the truth over all ocean grid points that are covered by sea ice in the neXtSIM-OPA simulation (concentration >0.15>0.15). To estimate a normalised RMSE, we normalise the forecast errors by the standard deviation of the climatology. The spread-skill ratio is estimated as ratio between the root-mean-squared standard deviation from GenSIM’s ensemble members and the RMSE of the ensemble mean with a correction factor of $\sqrt{17/16}$ (?). The energy spectra are estimated over the central Arctic region with $128\times 128$ grid points and averaged in logarithmic space over all initialisation times. To estimate the extent accuracy, the predicted sea-ice concentration is converted into an extent mask (concentration >0.15>0.15). The accuracy is then the agreement between the prediction and the truth over all ocean grid points.
Using the root mean-squared error (RMSE) as metric, GenSIM’s ensemble mean outperforms these baselines (Fig. S2a). Across all six predicted variables, GenSIM improves the scores, reaching an improvement of around $20\,\%$ compared to the deterministic forecasts. In the central Arctic, which is mostly year-long covered by sea ice, GenSIM achieves substantial improvements, and remarkably also in regions with marginal ice zones (MIZ), though small deficiencies in the sea-ice concentration during the change from melting to refreezing remains (see also Fig. S3 and Fig. S4). These results are striking: while the deterministic model is specifically trained to minimise the RMSE, GenSIM–designed as generative flow–achieves an indirect reduction. Derived from independent predictions, GenSIM’s ensemble mean is a true statistical average, explaining why it can outperform the deterministic model (?). Although well-tuned for 12 hours, GenSIM’s ensemble becomes seemingly underdispersive with increasing lead time, leading to typical spread-skill ratios of $0.8$ after 10 days. This underdispersion could be remedied by injecting more stochasticity (?) or tuning on the validation dataset (?) from which we refrain.
Like the geophysical model neXtSIM, GenSIM retains the energy at all scales even after a 10-day forecast (Fig. S2c), while the ensemble mean forecast and the deterministic model lose small-scale energy. This loss of small-scale information effectively smooths the forecast out (?, ?), leading to physical inconsistencies (?) and even instabilities in long-term simulations. GenSIM additionally predicts the sea-ice extent more accurately, which becomes especially evident during summer and autumn (Fig. S2d). GenSIM reprents better the sea-ice edge and marginal ice zone (MIZ), which we partially attribute to the generative nature of the model and partially to the novel censored loss function that explicitly incorporates the physical bounding of the variables. In conclusion, GenSIM outperforms the deterministic model and resolves the problems with the loss of small-scale information, like similarly found for the atmosphere (?), while better constraining the MIZ.
Score cards for short-term predictions
@@
Inspired by ECMWF’s score cards (https://sites.ecmwf.int/ifs/scorecards/, last access: 19th August 2025), the root-mean-squared error is decomposed into averages across seasons and regions. For that decomposition, we use the indexed regions as indicated in Fig. S1. Mesh cells over the Atlantic ocean that mostly cover open water are omitted from the estimation of the root-mean-squared error (RMSE). As previously, the RMSE is estimated only over ice-covered mesh cells (concentration >0.15>0.15). The score cards (Fig. S3 and Fig. S4) show how GenSIM’s ensemble mean improves the predictions compared to the deterministic surrogate model. Note that there is almost no snow during summer (JAS), making its error estimate unreliable in that season.